# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v2

@@ Abstract
Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex, the corresponding user questions also grow more challenging, leading PLMs with limited comprehension capabilities to produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods for PLMs, which, in turn, restricts the applications of PLM-based systems. Most recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale remains increasing. Therefore, integrating the LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
@@ I Introduction
Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23]. Incremental research on PLM-based optimization, such as table content encoding [24, 19, 25] and pre-training [20, 26], has further advanced this field. Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
@@
[…]
LLM-based Text-to-SQL
Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37] Spider-$S$ [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42] SParC [43] Knowledge- augmented BIRD [33], SQUALL [44], Spider-DK [39] Cross- lingual DuSQL [34], CSpider [42] Context- dependent CoSQL [35], Spider-$S$ [38], Spider-CG [38], SparC [43] Robustness ADVETA [37], Spider-SYN [40], Spider-Realistic [41] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [13], Exact Matching (EM) [13] Execution based Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33] Methods IV) In-context Learning Paradigm IV-A) Zero-shot [7], [33], [27], [45], [46], [47], [48], [49], [9], [50] Few-shot [33], [8], [51], [52], [53], [54], [49], [9], [32], [55] Decomposition Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58], SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61] Prompt Optimization DESEM+P [62], StructGPT [63], SD+SA+Voting [52]RAG+SP&DRC [64], C3 [30], DAIL-SQL [9] ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31] Reasoning Enhancement CoT [51, 33, 9, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55] SQL-CRAFT [55], FUXI [66] Execution Refinement MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62] DIN-SQL [8], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57] DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61] Fine-tuning Paradigm IV-B) Supervised Fine-tuning [45], [9], [50], [53] Enhanced Architecture CLLMs [69] Pre-training CodeS [10] Data Augmentation DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70] Decomposition DTS-SQL [71]
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
@@ IV-A In-context Learning
Methods Applied LLMs Benchmark Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
@@
As pilot experiments, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], achieve better performance rather than zero-shot approaches. [33] provides a 1-shot example combining with CoT [102] prompting, the decomposed reasoning steps of the given example trigger the text-to-SQL model for generating accurate SQL. [55] study the effect of the number of few-shot examples. [52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison. Furthermore, above the similarity selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples. A study of difficult-level samples selection [51] compared the performance of few-shot Codex, with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41]. Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels. [49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting. In their evaluations, they also test the impact of different input schema styles and different static and dynamic exemplar numbers.
The impact of cross-domain few-shot examples is also being studied [54]. When incorporating in-domain and out-of-domain with different numbers of examples, the in-domain demonstration always outperforms zero-shot and out-of-domain examples and gets better as the number of examples rises. To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches. The former style splits the schema, the column names, and the primary and foreign keys by bar, and the latter organizes them as natural language descriptions.
@@
As an intuitive solution, decomposing a challenging user question into simpler sub-questions and using multi-step reasoning for implementation can reduce the complexity of the full text-to-SQL task [51, 8]. Dealing with less complexity, LLMs have the potential to have better performance. The decomposition approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposing, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into smaller effective sub-tasks (e.g., schema linking, domain classification). (2) sub-question decomposing, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query.
@@
DIN-SQL [8] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction. DIN-SQL first finishes the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification. Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL. This approach comprehensively considers the decomposition of both sub-tasks and sub-questions. Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction. Refer to the Chain-of-Thought [102] and Least-to-Most prompting [103], QDecomp [51] introduce question decomposition prompting, which follows the question reduction stage in least-to-most prompting and instruct the LLM to decompose the original complex question as the intermediate reasoning steps. C3 [30] consists of three key components: clear prompting, calibration bias prompting, and consistency; these components are accomplished by assigning ChatGPT with different tasks. Firstly, the clear prompting component generates the schema linking and the distilled question-relevant schema as a clear prompt. Then, a multi-turn dialogue about text-to-SQL hints is utilized as a calibration bias prompt, which combines with the clear prompt to guide the SQL generation. The generated SQL queries are selected by consistency and execution-based voting to get the final SQL. MAC-SQL [57] presents a multi-agent collaborating framework; the text-to-SQL process is finished as the collaboration of the agents: Selector, Decomposer, and Refiner. The Selector preserves relevant tables for user questions; the Decomposer breaks down user questions into sub-questions and provides solutions; finally, the Refiner validates and refines the defective SQL. DEA-SQL [58] introduces a workflow paradigm aiming to enhance the attention and problem-solving scope of LLM-based text-to-SQL through decomposition. This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks. Through the workflow paradigm, an accurate SQL query is generated. SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation. Specifically, the framework constructs a graph structure for the user question and the corresponding database respectively, then uses the encoded graphs to construct structure linking [104, 105]. A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL. MetaSQL [59] introduces a three-stage approach for SQL generation, which consists of decomposition, generation, and rank. The decomposition stage uses semantic decomposition and metadata composition to process the user question. Taking the previously processed data as input, a text-to-SQL model using metadata-conditioned generation to generate some candidate SQL queries. Finally, a two-stage ranking pipeline is applied to get a global-optimal SQL query. PET-SQL [60] proposed a prompt-enhanced two-stage framework. Firstly, an elaborated prompt instructs the LLMs to generate preliminary SQL (PreSQL) where some few-shot demonstrations are selected based on similarity. Then, schema linking is found based on PreSQL and combined to prompt the LLMs to generate the Final SQL (FinSQL). Finally, multiple LLMs are utilized to generate a FinSQL, ensuring consistency based on the execution results.
@@
As previously introduced, few-shot learning is widely studied for prompting LLMs [84]. For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to achieve higher performance. Since the performance of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [106], many decisive factors that can influence the quality of the prompt have become focuses of the research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints). The process of prompt quality improvement is actually the prompt’s optimization, including few-shot sampling strategies, schema augmentation, and external knowledge generation.
@@
DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval. The framework first employs domain-specific words masking module to remove the semantic tokens in questions that preserve the question’s intentions. And then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation. The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names. With difficulty-based selection, the few-shot examples for QDecomp are difficult-level sampled. Besides similarity-diversity sampling, [52] proposed SD+SA+Voting (Similarity-Diversity+Schema Augmentation+Voting) sampling strategy. They first employ semantic similarity and $k$-Means cluster diversity for sampling few-shot examples and then enhance the prompt with schema knowledge (semantic or structure augmentation). C3 [30] framework comprises a clear prompting component, which takes the question and schema as the LLMs input, generates a clear prompt that includes a schema that removes the redundant information irrelevant to the user question and a schema linking, and also a calibration component providing hints. The LLMs take their composition as context-augmented prompts for SQL generation. A retrieval-augmented framework is introduced with sample-aware prompting [64], which simplifies the original question and extracts the question skeleton from the simplified question, then finishes the sample retrieval in the repository according to skeleton similarities. The retrieved samples are combined with the original question for few-shot prompting. ODIS [54] introduces the selection of a sample with out-of-domain demonstrations and in-domain synthetic data, which retrieves few-shot demonstrations from hybrid sources to augment the prompt representations. DAIL-SQL [9] proposed a novel approach to address the issues in few-shot sampling and organization, presenting a better balance between the quality and quantity of few-shot examples. DAIL Selection first masks domain-specific words in user and few-shot example questions, then ranks the candidate examples based on the embedded Euclidean distance. Meanwhile, the similarity between the pre-predicted SQL queries is calculated. Finally, the selection mechanism obtains the similarity-sorted candidates according to the pre-set criteria. The few-shot examples are guaranteed good similarity with both questions and SQL queries with this method. ACT-SQL [49] proposed dynamic examples in few-shot prompting, which is selected according to similarity score. FUSED [65] are presented to build a high-diversity demonstrations pool through human-free multiple-iteration synthesis to improve the diversity of the few-shot demonstrations. The pipeline of FUSED samples the demonstrations to be fused by clustering, then fuse the sampled demonstrations to construct the pool to enhance few-shot learning. Knowledge-to-SQL [31] framework aims to build a Data Expert LLM (DELLM) to provide knowledge for SQL generation. The DELLM is trained by supervised fine-tuning using human expert annotations [33] and further refined by preference learning with the database’s feedback. DELLM generates four categories of knowledge, the well-designed methods (e.g. DAIL-SQL [9], MAC-SQL [57]) incorporating the generated knowledge to achieve better performance for LLM-based text-to-SQL with in-context learning.
@@
LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [107]. Since for the text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41], the prompting strategies for the LLMs reasoning possess the potential to enhance their SQL generation capabilities. Recent studies primarily focus on incorporating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require multi-step reasoning and the issue of self-consistency [108] in SQL generation.
@@
To design criteria for accurate SQL generation, whether a generated SQL can be successfully executed and elicit a correct answer for the user question is always the priority [13]. As a complex programming task, generating the correct SQL in one go becomes challenging. Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30]. The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for regeneration, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining feedback from the database. This feedback might be an error, or it might yield results that will be appended to the second round prompt. Through in-context learning of this feedback, LLMs are able to refine or regenerate the original SQL, thereby enhancing accuracy. 2) Utilize execution-based selection strategies for generated SQL, sample multiple generated SQL queries from LLM), and execute each in the database. Based on the results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define a query from the SQL set that satisfies the criteria as the final predicted SQL.

### v3

@@ Abstract
Generating accurate SQL from natural language questions (text-to-SQL) is a long-standing challenge due to the complexities in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex, the corresponding user questions also grow more challenging, causing PLMs with parameter constraints to produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems. Recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale increases. Therefore, integrating LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
Index Terms:
text-to-SQL, large language models, database, natural language understanding
@@ I Introduction
Previous studies have made notable progress in implementing text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23]. Incremental research on PLM-based optimization, such as table content encoding [24, 19, 25] and pre-training [20, 26], has further advanced this field. Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [LLM-based Text-to-SQL, leaf-head, ver [ Datasets
III-A), leaf-datasets, text width=5em [ Original
Datasets
Post-annotated
Datasets, leaf-datasets, text width=6em [ Cross-
domain, leaf-datasets, text width=5.5em [BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37],
Spider-SS [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42],
SParC [43], modelnode-datasets, text width=43.5em] ] [ Knowledge-
augmented, leaf-datasets, text width=5.5em [BIRD [33], SQUALL [44], Spider-DK [39], modelnode-datasets, text width=43.5em] ] [ Cross-
lingual, leaf-datasets, text width=5.5em [DuSQL [34], CSpider [42], modelnode-datasets, text width=43.5em] ] [ Context-
dependent, leaf-datasets, text width=5.5em [CoSQL [35], Spider-SS [38], Spider-CG [38], SparC [43], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [37], Spider-SYN [40], Spider-Realistic [41], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
Metrics
III-B), leaf-metrics,text width=5em [ Content-
Matching
based, leaf-metrics, text width=6em [Component Matching (CM) [13], Exact Matching (EM) [13], modelnode-metrics, text width=50.7em] ] [ Execution
based, leaf-metrics, text width=6em [Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33], modelnode-metrics, text width=50.7em] ] ] [ Methods
IV), leaf-methods,text width=5em [ In-context
Learning
Paradigm
IV-A), leaf-methods, text width=6em [ Trivial
Prompt, leaf-methods, text width=6.5em [ Zero-shot, modelnode-methods, text width=4.5em [[7], [33], [27], [45], [46], [47], [48], [49], [9], [50], modelnode-methods,text width=36.2em]] [ Few-shot, modelnode-methods, text width=4.5em [[33], [8], [51], [52], [53], [54], [49], [9], [32], [55] , modelnode-methods,text width=36.2em]] ] [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58],
SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] [ Prompt
Optimization, leaf-methods, text width=6.5em [DESEM+P [62], StructGPT [63], SD+SA+Voting [52], RAG+SP&DRC [64], C3 [30], DAIL-SQL [9],
ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31], modelnode-methods,text width=42.5em] ] [ Reasoning
Enhancement, leaf-methods, text width=6.5em[CoT [51, 33, 9, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55],
SQL-CRAFT [55], FUXI [66], modelnode-methods,text width=42.5em] ] [ Execution
Refinement, leaf-methods, text width=6.5em[MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62],
DIN-SQL [8], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57],
DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
Paradigm
IV-B), leaf-methods,text width=6em [ Vanilla
Fine-tuning, leaf-methods, text width=6.5em [[45], [9], [50], [53], modelnode-methods,text width=42.5em] ] [ Enhanced
Architecture, leaf-methods, text width=6.5em [CLLMs [69], modelnode-methods,text width=42.5em] ] [ Pre-training, leaf-methods, text width=6.5em [CodeS [10], modelnode-methods,text width=42.5em] ] [ Data
Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70], modelnode-methods,text width=42.5em] ] [ Decomposition, leaf-methods, text width=6.5em [DTS-SQL [71], modelnode-methods,text width=42.5em] ] ] ] ]
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. Related works in recent years have surveyed the text-to-SQL systems in deep learning approaches and provided insights into the previous deep neural network and PLM-based research [74, 2, 29]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
@@ IV-A In-context Learning
Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
@@
As empirical studies, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], exhibit a solid performance during the comparison with zero-shot prompting. [33] provides a 1-shot detailed example to trigger the text-to-SQL model for generating accurate SQL. [55] study the effect of the number of few-shot examples. [52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison. Furthermore, above the similarity-based selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples. A study of difficult-level samples selection [51] compared the performance of few-shot Codex [100], with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41]. Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels. [49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting. In their settings, they also evaluate the impact of different input schema styles and various static and dynamic exemplar numbers.
The impact of cross-domain few-shot examples is also being studied [54]. When incorporating in-domain and out-of-domain examples with different numbers, the in-domain demonstration outperforms zero-shot and out-of-domain examples, which will also perform better as the number of examples rises. To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches. The former style splits the schema, the column names, and the primary and foreign keys by bar, and the latter organizes them as natural language descriptions.
@@
As an intuitive solution, decomposing a challenging user question into simpler sub-questions or using multi-components for implementation can reduce the complexity of the overall text-to-SQL task [51, 8]. Dealing with less complexity, LLMs have the potential to elicit more accurate SQL generation. The decomposition-based approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposition, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into more manageable effective sub-tasks (e.g., schema linking [71], domain classification [54]). (2) sub-question decomposition, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query. The technical novelty of the decomposition paradigm
@@
DIN-SQL [8] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction. DIN-SQL first generates the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification. Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL. This approach comprehensively considers the sub-questions decomposition as a module of sub-tasks decomposition. Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction. Refer to the Chain-of-Thought [103] and Least-to-Most prompting [104], QDecomp [51] introduce question decomposition prompting, which follows the question reduction stage in least-to-most prompting and instruct the LLM to decompose the original complex question as the intermediate reasoning steps. C3 [30] consists of three key components: clear prompting, calibration bias prompting, and consistency; these components are accomplished by assigning ChatGPT with different tasks. Firstly, the clear prompting component generates the schema linking and the distilled question-relevant schema as a clear prompt. Then, a multi-turn dialogue about text-to-SQL hints is utilized as a calibration bias prompt, which combines with the clear prompt to guide the SQL generation. The generated SQL queries are selected by consistency and execution-based voting to get the final SQL. MAC-SQL [57] presents a multi-agent collaborating framework; the text-to-SQL process is finished as the collaboration of the agents: Selector, Decomposer, and Refiner. The Selector preserves relevant tables for user questions; the Decomposer breaks down user questions into sub-questions and provides solutions; finally, the Refiner validates and refines the defective SQL. DEA-SQL [58] introduces a workflow paradigm aiming to enhance the attention and problem-solving scope of LLM-based text-to-SQL through decomposition. This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks. The workflow paradigm allows the LLM to generate more accurate SQL queries. SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation. Specifically, the framework constructs a graph structure for the user question and the corresponding database respectively, then uses the encoded graphs to construct structure linking [105, 106]. A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL. MetaSQL [59] introduces a three-stage approach for SQL generation: decomposition, generation, and ranking. The decomposition stage uses semantic decomposition and metadata composition to process the user question. Taking the previously processed data as input, a text-to-SQL model using metadata-conditioned generation to generate some candidate SQL queries. Finally, a two-stage ranking pipeline is applied to get a global-optimal SQL query. PET-SQL [60] proposed a prompt-enhanced two-stage framework. Firstly, an elaborated prompt instructs the LLMs to generate preliminary SQL (PreSQL) where some few-shot demonstrations are selected based on similarity. Then, schema linking is found based on PreSQL and combined to prompt the LLMs to generate the Final SQL (FinSQL). Finally, multiple LLMs are utilized to generate a FinSQL, ensuring consistency based on the execution results.
@@
As previously introduced, few-shot learning is widely studied for prompting LLMs [85]. For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to elicit better performance. Since the accuracy of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [107], many decisive factors that can influence the quality of the prompt have become points of focus of the current research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints) The process of improving prompt quality can be summarized as the prompt’s optimization, including advanced few-shot sampling strategies, schema information augmentation, and external knowledge integration.
@@
DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval. The framework first employs domain-specific words masking module to remove the semantic tokens in user questions that preserve the intentions. It then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation. The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names. With difficulty-based selection, the few-shot examples for QDecomp are difficult-level sampled. Besides similarity-diversity sampling, [52] proposed SD+SA+Voting (Similarity-Diversity+Schema Augmentation+Voting) sampling strategy. They first employ semantic similarity and $k$-Means cluster diversity for sampling few-shot examples and then enhance the prompt with schema knowledge (semantic or structure augmentation). C3 [30] framework comprises a clear prompting component, which takes the question and schema as the LLMs input, generates a clear prompt that includes a schema that removes the redundant information irrelevant to the user question and a schema linking, and also a calibration component providing hints. The LLMs take their composition as context-augmented prompts for SQL generation. A retrieval-augmented framework is introduced with sample-aware prompting [64], which simplifies the original question and extracts the question skeleton from the simplified question, then finishes the sample retrieval in the repository according to skeleton similarities. The retrieved samples are combined with the original question for few-shot prompting. ODIS [54] introduces the selection of a sample with out-of-domain demonstrations and in-domain synthetic data, which retrieves few-shot demonstrations from hybrid sources to augment the prompt representations. DAIL-SQL [9] proposed a novel approach to address the issues in few-shot sampling and organization, presenting a better balance between the quality and quantity of few-shot examples. DAIL Selection first masks domain-specific words in user and few-shot example questions, then ranks the candidate examples based on the embedded Euclidean distance. Meanwhile, the similarity between the pre-predicted SQL queries is calculated. Finally, the selection mechanism obtains the similarity-sorted candidates according to the pre-set criteria. The few-shot examples are guaranteed good similarity with both questions and SQL queries with this method. ACT-SQL [49] proposed dynamic examples in few-shot prompting, which is selected according to similarity score. FUSED [65] are presented to build a high-diversity demonstrations pool through human-free multiple-iteration synthesis to improve the diversity of the few-shot demonstrations. The pipeline of FUSED samples the demonstrations to be fused by clustering, then fuse the sampled demonstrations to construct the pool to enhance few-shot learning. Knowledge-to-SQL [31] framework aims to build a Data Expert LLM (DELLM) to provide knowledge for SQL generation. The DELLM is trained by supervised fine-tuning using human expert annotations [33] and further refined by preference learning with the database’s feedback. DELLM generates four categories of knowledge, the well-designed methods (e.g. DAIL-SQL [9], MAC-SQL [57]) incorporating the generated knowledge to achieve better performance for LLM-based text-to-SQL with in-context learning.
@@
LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [108]. For text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41]. The prompting strategies for LLMs’ reasoning have the potential to enhance SQL generation capabilities. Recent studies primarily focus on integrating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require sophisticated reasoning [^3] The difference between multi-step reasoning (e.g., Chain of Thought) and the decomposition paradigm is that the former research focuses on advancing inherent reasoning within a single-turn generation, whereas the latter study involves using different components to assist the final generation through multiple calls of LLMs. and the self-consistency in SQL generation.
@@
When designing criteria for accurate SQL generation, the priority is always whether a generated SQL can be successfully executed and retrieve content to correctly answer the user question [13]. As a complex programming task, generating the correct SQL in one go is challenging. Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30]. The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for re-generation, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining the database’s feedback. This feedback might be an error, or it might yield results that will be appended to the second round prompt. Through in-context learning of this feedback, LLMs are able to refine or re-generate the original SQL, thereby enhancing accuracy. 2) Utilize execution-based selection strategies for the generated SQL, sample multiple generated SQL queries from LLM, and execute each in the database. Based on the execution results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define one SQL query from the SQL set that satisfies the criteria as the final predicted SQL.